# MetaSym: A Symplectic Meta-learning Framework for Physical Intelligence

### v2

@@ Appendix B Experimental Setup
B.1 Spring Mesh System
Spring networks are a simple proxy for numerous physical scenarios involving deformable solids and cloth (in computer graphics, mechanics, or robotics). Each pair of connected particles exchanges spring forces that depend on displacements from rest lengths. Viscous damping terms further shape the evolution. While the individual dynamics (Hooke’s law) are straightforward, the combination of hundreds of coupled springs can give rise to complex large-scale deformations and oscillations. By benchmarking on a spring-mesh, we can examine how MetaSym learns large deformations, wave propagation through a membrane, or the impact of damping.
The training dataset consists of 25 distinct spring-mesh systems, each characterized by a unique set of physical parameters, including spring stiffness, damping coefficients, and initial conditions as indicated by Table 2(a). These parameters are sampled from a predefined distribution to ensure sufficient diversity within the training set. Each system is simulated over a time span of 2000 irregular timesteps, capturing the full trajectory of node displacements and momenta. The resulting dataset provides a rich representation of dynamical behaviors within the parameter space.
To assess generalization and robustness, we construct a test dataset comprising ten additional spring-mesh systems. Unlike the training set, the parameters for these systems are drawn from distributions that differ from those used during training as indicated by Table 2(b), introducing a domain shift that mimics real-world variations. This OOD test set enables a rigorous evaluation of the model’s ability to extrapolate beyond the observed training dynamics and adapt to unseen conditions. For further information regarding the spring-mesh benchmark refer to [1].
@@ B.1 Spring Mesh System
Mesh Size [$n_{x}\times n_{y}$] 10 $\times$ 10
(a) In-distribution parameters $T=2000$
@@
Mesh Size [$n_{x}\times n_{y}$] 10 $\times$ 10
(b) Out-of-distribution parameters $T=2000$
@@ B.2 Open Quantum System Derivation
Understanding the behavior of quantum systems plays a vital role in the development of promising technologies such as quantum computing, metrology, and sensing. While deep learning has found great success in several areas such as quantum control [45], error correction [47] and tuning [48, 49], predicting the measurement record based on modeling quantum dynamics has long remained elusive. In many scenarios, the system of interest is open: it couples to an environment (or bath) that can introduce thermal noise and dissipation. Furthermore, continuous monitoring (e.g., via homodyne detection) adds additional measurement backaction, reflecting fundamental constraints from quantum mechanics [50]. Capturing these noise and measurement effects is pivotal for accurately predicting quantum trajectories and devising robust control protocols.
Unlike closed Hamiltonian evolutions, open quantum systems require one to solve Stochastic Master Equations (SMEs) incorporating decoherence and measurement terms. These equations produce trajectories of the (mixed) quantum state conditioned on the noisy measurement record. In many practical settings, however, we only have direct access to certain observables (e.g., position and momentum quadratures) rather than the full quantum state. Hence, training a deep learning network to model the quantum system and to predict future measurement outcomes becomes a natural and practically relevant challenge. The SME describes the evolution of the conditioned quantum state $\rho_{c}(t)$ under the effect of environmental and measurement noise as [50]
@@
where the dissipator is $\mathcal{D}[\hat{L}]\rho=\hat{L}\rho\hat{L}^{\dagger}-\frac{1}{2}\left\{\hat{L}^{\dagger}\hat{L},\rho\right\}$. Each Linblad operator, $\hat{L}_{j}$, represents a collapse operator that encodes coupling to the environment (e.g., photon loss to a reservoir, thermal excitations, dephasing, etc.) [51]. The stochastic backaction term, $\mathcal{H}[\hat{M}]\rho=\hat{M}\rho+\rho\hat{M}^{\dagger}-\operatorname{Tr}\left[\left(\hat{M}+\hat{M}^{\dagger}\right)\rho\right]\rho$, describes continuous monitoring of an observable $\hat{M}$, with $\eta$ the measurement efficiency ($0\leq\eta\leq 1$). The Wiener increment, $dW_{t}$, captures the randomness inherent in quantum measurement outcomes.
@@ B.2.1 Setup for Parameteric Oscillator
with $\omega$, $\chi$ and $\beta$ being the oscillator frequency, squeezing strength and cubic driving term respectively. We also include two Linblad operators:
@@
A common measurement technique often employed in experimental settings is called heterodyne measurement. Heterodyne detection continuously measures both quadratures of the output of our dissipative qubit by mixing it with a local oscillator at a slightly shifted frequency and then demodulating the resulting beat signal. This yields two simultaneous photocurrents, often referred to as the in-phase ($I$) and quadrature ($Q$) components. This reflects a key practical outcome of solving Eq. (6) is that $\rho_{c}(t)$ depends on this random measurement trajectory.
We employ heterodyne detection of the field operator $\hat{a}$. Based on this measurement scheme, we can get seperate the real and imaginary parts to obtain quadrature values that roughly correspond to $X$ and $P$ while adding quantum noise and measurement uncertainities due to quantum mechanical effect [46].
In the following Tables 3(b) & 3(a), we describe the parameters used to generate our dataset by solving the SME in order to generate training data.
@@ B.3 Quadrotor
The quadrotor system challenges data-driven methods with its floating-base dynamics. To generate the training and OOD validation datasets we use Crocoddyl trajectory optimization package based on the dynamics model proposed by [52]. The training dataset comprises 30 systems with randomized parameters such as inertia, torque constant and rotor lengths as indicated in Table 4(a). Each trajectory is generated from a randomized initial condition to a random terminal position with zero velocity, within a pre-set bounding box, to avoid unrealistic velocities.
@@ C.1 Overall Map
We aim to show that if $\Phi_{\theta_{AD}}$ remains a small (bounded) perturbation $O(\rho)$ from identity in $\mathbf{z}_{c}$-space, then $\Phi_{\theta}$ preserves the symplectic structure up to a small error term. This is called $O(\rho)$ near-symplecticity. We assume that the dissipation $\mathbf{d}_{t}$ and control-input $\mathbf{u}_{t}$ are seperable for the sake of this proof.
@@ C.2 SymplecticEncoder
From Jin et al., we know that LA-SympNets are fully symplectic due to their construction [12]. By extension, we can show that our SymplecticEncoder $\Phi_{{\theta}_{SE}}$ is symplectic.
@@
$J\;=\;\begin{pmatrix}0&I_{d}\\[3.0pt] -I_{d}&0\end{pmatrix},$
@@
$\mathrm{d}\Psi(\mathbf{x})^{\top}\,J\,\mathrm{d}\Psi(\mathbf{x})\;=\;J\quad\forall\,\mathbf{x}\in\mathcal{X}.$
@@
$\mathrm{d}\Phi_{\theta_{SE}}(\mathbf{x})^{\top}\,J\,\mathrm{d}\Phi_{\theta_{SE}}(\mathbf{x})=J,\quad\det(\mathrm{d}\Phi_{\theta_{SE}}(\mathbf{x}))=1.$
@@ C.3 Decoder with Control and Dissipation
  • $\mathcal{Z}\subseteq\mathbb{R}^{2d}$ is the latent phase-space output by $\Phi_{\theta_{SE}}$,
@@
  • $\mathcal{D}\subseteq\mathbb{R}^{r}$ represents dissipative parameters (bounded by $\|\mathbf{d}_{t}\|\leq D_{\max}$), which model forces that remove or drain energy (e.g. friction or drag).
The decoder modifies the latent state by
@@
where $F_{\theta_{AD}}$ can be cross-attention with magnitude modulated by $\|\mathbf{u}_{t}\|$, or a damping formula modulated by $\|\mathbf{d}_{t}\|$.
@@ C.4 $O(\rho)$ Near-Symplectic Proof and Explicit Bound
In this section, we prove that if $\mathbf{z}_{c}\mapsto\mathbf{z}_{c}+F_{\theta_{AD}}(\mathbf{z}_{c})$ is a bounded perturbation from identity (in partial derivatives), then the composition with a strict symplectic map remains close to preserving $J$.
@@
for some constants $\alpha_{0},\alpha_{u},\alpha_{d}\geq 0$. This covers cross-attention (scaled by $\|\mathbf{u}_{t}\|$) and dissipative (scaled by $\|\mathbf{d}_{t}\|$) terms. Since $\|\mathbf{u}_{t}\|\leq U_{\max}$ and $\|\mathbf{d}_{t}\|\leq D_{\max}$, we define:
@@
Hence, $\max_{(\mathbf{z}_{c},\mathbf{u}_{t},\mathbf{d}_{t})}\bigl{\|}\mathrm{d}_{\mathbf{z}_{c}}\Phi_{\theta_{AD}}-I\bigr{\|}\leq\rho$. Equivalently,
@@
$\mathrm{d}\Phi_{\theta}(\mathbf{x},\mathbf{u}_{t},\mathbf{d}_{t})=\underbrace{\mathrm{d}_{\mathbf{z}_{c}}\!\Phi_{\theta_{AD}}\bigl{(}\mathbf{z}_{c},\mathbf{u}_{t},\mathbf{d}_{t}\bigr{)}}_{\textstyle I+A,\;\|A\|\leq\rho}\quad\times\underbrace{\mathrm{d}\Phi_{\theta_{SE}}(\mathbf{x})}_{\textstyle\text{strictly symplectic.}}$
@@
$\mathrm{d}\Phi_{\theta_{SE}}(\mathbf{x})^{\top}\;J\;\mathrm{d}\Phi_{\theta_{SE}}(\mathbf{x})=\;J\quad\text{and}\quad\det(\mathrm{d}\Phi_{\theta_{SE}}(\mathbf{x}))=1.$
@@
$\bigl{\|}\mathrm{d}\Phi_{\theta}(\mathbf{x})^{\top}\,J\,\mathrm{d}\Phi_{\theta}(\mathbf{x})-J\bigr{\|}\;\leq\;C\,\rho,$
@@
$\det\bigl{(}\mathrm{d}\Phi_{\theta}(\mathbf{x})\bigr{)}=\mathbf{1}+O(\rho),$
@@
Let $\mathbf{x}=(\mathbf{q}_{t},\mathbf{p}_{t})$, $\mathbf{z}_{c}=\Phi_{\theta_{SE}}(\mathbf{x})$, and $I+A=\mathrm{d}_{\mathbf{z}_{c}}\!\Phi_{\theta_{AD}}(\mathbf{z}_{c},\mathbf{u}_{t},\mathbf{d}_{t})$ with $\|A\|\leq\rho$. Then
$\mathrm{d}\Phi_{\theta}(\mathbf{x},\mathbf{u}_{t},\mathbf{d}_{t})=(I+A)\,\mathrm{d}\Phi_{\theta_{SE}}(\mathbf{x}).$
@@
$\mathrm{d}\Phi_{\theta}^{\top}\,J\,\mathrm{d}\Phi_{\theta}=\mathrm{d}\Phi_{\theta_{SE}}^{\top}\,(I+A)^{\top}\,J\,(I+A)\,\mathrm{d}\Phi_{\theta_{SE}}.$
Expanding $(I+A)^{\top}J(I+A)=J+A^{\top}J+JA+A^{\top}JA=J+O(\|A\|)$, we substitute $\mathrm{d}\Phi_{\theta_{SE}}^{\top}\,J\,\mathrm{d}\Phi_{\theta_{SE}}=J$:
$\mathrm{d}\Phi_{\theta}^{\top}\,J\,\mathrm{d}\Phi_{\theta}=J+O(\|A\|)=J+O(\rho).$
In operator norm, $\|\mathrm{d}\Phi_{\theta}^{\top}\,J\,\mathrm{d}\Phi_{\theta}-J\|\leq C\,\rho$. Since $\mathrm{d}\Phi_{\theta_{SE}}$ is volume-preserving, $\det(\mathrm{d}\Phi_{\theta})=\det(I+A)\times\mathbf{1}=\mathbf{1}+O(\rho)$. ∎
@@
$F_{\theta_{AD}}(\mathbf{z}_{c},\mathbf{u}_{t},\mathbf{d}_{t})=\;\alpha\,\mathrm{CrossAttn}(\mathbf{z}_{c},\mathbf{u}_{t})\;-\;\gamma\,(\mathbf{d}_{t}^{\top}\!\!\star\mathbf{z}_{c}),$
where $\mathbf{d}_{t}^{\top}\!\!\star\mathbf{z}_{c}$ indicates some parametric dissipator. If $\mathrm{CrossAttn}$ has partial derivative in $\mathbf{z}_{c}$ normed by $\|\mathbf{u}_{t}\|$, and $\mathbf{d}_{t}^{\top}\!\!\star\mathbf{z}_{c}$ is linear in $\|\mathbf{d}_{t}\|$, then
@@
We choose not to explicitly bound this perturbation during training to allow for MetaSym to model extremely dissipative systems. In Appendix D.2, we provide the perturbation bound $C\rho$ emperically for systems undergoing extreme dissipation and control. We observe that this bound is $<1$ even under these extreme settings.

### v3

@@ Appendix B Experimental Setup
B.1 Spring-mesh System
Spring networks are a simple proxy for numerous physical scenarios involving deformable solids and cloth (in computer graphics, mechanics, or robotics). Each pair of connected particles exchanges spring forces that depend on displacements from rest lengths. Viscous-damping terms further shape the evolution. While the individual dynamics (Hooke’s law) are straightforward, the combination of hundreds of coupled springs can give rise to complex large-scale deformations and oscillations. By benchmarking on a spring-mesh, we can examine how MetaSym learns large deformations, wave propagation through a membrane, or the impact of damping.
The training dataset consists of $25$ distinct spring-mesh systems, each characterized by a unique set of physical parameters, including spring stiffness, damping coefficients, and initial conditions as indicated by Table 2(a). These parameters are sampled from a predefined distribution to ensure sufficient diversity within the training set. Each system is simulated over a time span of $2000$ irregular time-steps, capturing the full trajectory of node displacements and momenta. The resulting dataset provides a rich representation of dynamical behaviors within the parameter space.
To assess generalization and robustness, we construct a test dataset comprising ten additional spring-mesh systems. Unlike the training set, the parameters for these systems are drawn from distributions that differ from those used during training as indicated by Table 2(b), introducing a domain shift that mimics real-world variations. This OOD test set enables a rigorous evaluation of the model’s ability to extrapolate beyond the observed training dynamics and adapt to unseen conditions. For further information regarding the spring-mesh benchmark refer to Otness et al. (2021).
@@ B.1 Spring-mesh System
Mesh Size $n_{x}\times n_{y}$] 10 $\times$ 10
(b) Out-of-distribution parameters $T=2000$
@@
Mesh Size $n_{x}\times n_{y}$] 10 $\times$ 10
@@ B.2 Open Quantum System Derivation
Understanding the behavior of quantum systems plays a vital role in the development of promising technologies such as quantum computing, metrology, and sensing. While deep-learning has found great success in several areas such as quantum control (Vaidhyanathan et al., 2024), error correction (Bausch et al., 2024) and tuning (Ares, 2021; Gebhart et al., 2023), predicting the measurement record based on modeling quantum dynamics has long remained elusive. In many scenarios, the system of interest is open: it couples to an environment (or bath) that can introduce thermal noise and dissipation. Furthermore, continuous monitoring (e.g., via homodyne detection) adds additional measurement backaction, reflecting fundamental constraints from quantum mechanics (Jacobs & Steck, 2006). Capturing these noise and measurement effects is pivotal for accurately predicting quantum trajectories and devising robust control protocols.
Unlike closed Hamiltonian evolutions, open quantum systems require one to solve Stochastic Master Equations (SMEs) incorporating decoherence and measurement terms. These equations produce trajectories of the (mixed) quantum state conditioned on the noisy measurement record. In many practical settings, however, we only have direct access to certain observables (e.g., position and momentum quadratures) rather than the full quantum state. Hence, training a deep learning network to model the quantum system and to predict future measurement outcomes becomes a natural and practically relevant challenge. The SME describes the evolution of the conditioned quantum state $\rho_{c}(t)$ under the effect of environmental and measurement noise as in Jacobs & Steck (2006)
@@
where the dissipator is $\mathcal{D}[\hat{L}]\rho=\hat{L}\rho\hat{L}^{\dagger}-\frac{1}{2}\left\{\hat{L}^{\dagger}\hat{L},\rho\right\}$. Each Linblad operator, $\hat{L}_{j}$, represents a collapse operator that encodes coupling to the environment (e.g., photon loss to a reservoir, thermal excitations, dephasing, etc.) (Manzano, 2020). The stochastic backaction term, $\mathcal{H}[\hat{M}]\rho=\hat{M}\rho+\rho\hat{M}^{\dagger}-\operatorname{Tr}\left[\left(\hat{M}+\hat{M}^{\dagger}\right)\rho\right]\rho$, describes continuous monitoring of an observable $\hat{M}$, with $\eta$ the measurement efficiency ($0\leq\eta\leq 1$). The Wiener increment, $dW_{t}$, captures the randomness inherent in quantum measurement outcomes.
@@ B.2.1 Setup for Parameteric Oscillator
with $\omega$, $\chi$ and $\beta$ being the oscillator frequency, squeezing strength and cubic driving term respectively. We also include two Linblad operators,
@@
Figure 5: Quantum System’s position quadrature $X$ evolution: The faint gray traces represent individual single-shot trajectories, showing the inherent randomness caused by measurement backaction and thermal noise, directly visualizing the noisy nature of the dataset. The blue shaded region marks the confidence interval, quantifying the standard deviation of the quantum state’s fluctuations over time. This demonstrates the jagged unpredictability of individual runs. The ensemble average (solid blue line) converges to the smooth, and deterministic prediction of the Master Equation (red dashed line) over multiple trajectories.
A common measurement technique often employed in experimental settings is called heterodyne measurement. Heterodyne detection continuously measures both quadratures of the output of our dissipative system by mixing it with a local oscillator at a slightly shifted frequency and then demodulating the resulting beat signal. This yields two simultaneous photocurrents, often referred to as the in-phase ($I$) and quadrature ($Q$) components. This reflects a key practical outcome of solving Eq. 7 is that $\rho_{c}(t)$ depends on this random measurement trajectory.
@@
We employ heterodyne detection of the field operator $\hat{a}$. Based on this measurement scheme, we can get separate the real and imaginary parts to obtain quadrature values that roughly correspond to $X$ and $P$ while adding quantum noise and measurement uncertainties due to quantum mechanical effect (Flurin et al., 2020).
In the following Tables 3(a) & 3(b), we describe the parameters used to generate our dataset by solving the SME in order to generate training data.
We also include, Fig. 5, that illustrates the stochastic evolution of the position quadrature $X$ for our system. The nature of the system’s stochasticity is revealed by the shaded confidence interval, the shape of which is dictated by Eq. 7.
@@ B.3 Quadrotor
The quadrotor system challenges data-driven methods with its floating-base dynamics. To generate the training and OOD validation datasets we use Crocoddyl trajectory optimization package based on the dynamics model proposed by Geisert & Mansard (2016). The training dataset comprises $30$ systems with randomized parameters such as inertia, torque constant and rotor lengths as indicated in Table 4(a). Each trajectory is generated from a randomized initial condition to a random terminal position with zero velocity, within a pre-set bounding box, to avoid unrealistic velocities.
@@
The results in Fig. 6 verify the accuracy and effectiveness of our method for simulated floating-base dynamics systems. This consists the pre-training executed before fine-tuning the architecture on the real-world quadrotor dataset illustrated in the main text.
Figure 6: (Left) Represents the translational and angular phase-space evolution of the quadrotor, with training and test trajectories generated using the Crocoddyl trajectory optimization package (Mastalli et al., 2020). Each task is initialized with randomized initial conditions and a randomized terminal position over a $1.5s$ horizon ($150$ time-steps) with a context window of $30$ time-steps. The ground-truth test trajectory (orange line) overlaps with MetaSym’s predictions (blue line) indicating the excellent predictive capabilites of our model. (Right) Plots summarizing the mean squared error (MSE) of each of the phase-space coordinates evolution for five randomly generated test systems (dots). The consistently low median errors (boxes) across all components underscore the robustness and generalization capabilities of the model.
@@ C.1 Overall Map
We aim to show that if $\Phi_{\theta_{AD}}$ remains a small (bounded) perturbation $O(\rho)$ from identity in $\mathbf{z}_{c}$-space, then $\Phi_{\theta}$ preserves the symplectic structure up to a small error term. This is called $O(\rho)$ near-symplecticity. We assume that the dissipation $\mathbf{d}_{t}$ and control-input $\mathbf{u}_{t}$ are separable for the sake of this proof.
@@ C.2 SymplecticEncoder
From Jin et al. (2020), we know that LA-SympNets are fully symplectic due to their construction. By extension, we can show that our SymplecticEncoder $\Phi_{{\theta}_{SE}}$ is symplectic.
@@
$\mathbf{J}\;=\;\begin{pmatrix}\mathbf{0}&\mathbf{I}_{d}\\[3.0pt] -\mathbf{I}_{d}&\mathbf{0}\end{pmatrix}$,
@@
$\mathrm{d}\Psi(\mathbf{x})^{\top}\,\mathbf{J}\,\mathrm{d}\Psi(\mathbf{x})\;=\;\mathbf{J}\quad\forall\,\mathbf{x}\in\mathcal{X}.$
@@
$\mathrm{d}\Phi_{\theta_{SE}}(\mathbf{x})^{\top}\,\mathbf{J}\,\mathrm{d}\Phi_{\theta_{SE}}(\mathbf{x})=\mathbf{J},\quad\det(\mathrm{d}\Phi_{\theta_{SE}}(\mathbf{x}))=1.$
@@ C.3 Decoder with Control and Dissipation
  • $\mathcal{Z}\subseteq\mathbb{R}^{2d}$ is the latent phase-space output by ($\Phi_{\theta_{SE}}$),
@@
  • $\mathcal{D}\subseteq\mathbb{R}^{r}$ represents dissipative parameters (bounded by $\|\mathbf{d}_{t}\|\leq D_{\max}$), which model forces that remove or drain energy (e.g., friction or drag).
In order to account for the effects of the cross attention, the decoder modifies the latent state by
@@
where ($F_{\theta_{AD}}$) can be cross-attention with magnitude modulated by $\|\mathbf{u}_{t}\|$, or a damping formula modulated by $\|\mathbf{d}_{t}\|$.
@@ C.4 $O(\rho)$ Near-Symplectic Proof and Explicit Bound
In this section, we prove that if $\mathbf{z}_{c}\mapsto\mathbf{z}_{c}+F_{\theta_{AD}}(\mathbf{z}_{c})$ is a bounded perturbation from identity (in partial derivatives), then the composition with a strict symplectic map remains close to preserving $\mathbf{J}$.
@@
for some constants $\alpha_{0},\alpha_{u},\alpha_{d}\geq 0$. This covers cross-attention scaled by $\|\mathbf{u}_{t}\|$ and dissipative scaled by $\|\mathbf{d}_{t}\|$ terms. Since $\|\mathbf{u}_{t}\|\leq U_{\max}$ and $\|\mathbf{d}_{t}\|\leq D_{\max}$, we define:
@@
Hence, $\max_{\mathbf{z}_{c},\mathbf{u}_{t},\mathbf{d}_{t}}\bigl\|\mathrm{d}_{\mathbf{z}_{c}}\Phi_{\theta_{AD}}-\mathbf{I}\bigr\|\leq\rho$. Equivalently,
@@
$\mathrm{d}\Phi_{\theta}(\mathbf{x},\mathbf{u}_{t},\mathbf{d}_{t})=\underbrace{\mathrm{d}_{\mathbf{z}_{c}}\!\Phi_{\theta_{AD}}\bigl(\mathbf{z}_{c},\mathbf{u}_{t},\mathbf{d}_{t}\bigr)}_{\mathbf{I}+\mathbf{A},\;\|\mathbf{A}\|\leq\rho}\quad\times\underbrace{\mathrm{d}\Phi_{\theta_{SE}}(\mathbf{x})}_{\textstyle\text{strictly symplectic.}}$
@@
$\mathrm{d}\Phi_{\theta_{SE}}(\mathbf{x})^{\top}\;\mathbf{J}\;\mathrm{d}\Phi_{\theta_{SE}}(\mathbf{x})=\;\mathbf{J}\quad\text{and}\quad\det(\mathrm{d}\Phi_{\theta_{SE}}(\mathbf{x}))=1.$
@@
$\bigl\|\mathrm{d}\Phi_{\theta}(\mathbf{x})^{\top}\,\mathbf{J}\,\mathrm{d}\Phi_{\theta}(\mathbf{x})-\mathbf{J}\bigr\|\;\leq\;C\,\rho,$
@@
$\det\bigl(\mathrm{d}\Phi_{\theta}(\mathbf{x})\bigr)=1+O(\rho),$
@@
Let $\mathbf{x}=(\mathbf{q}_{t},\mathbf{p}_{t})$, $\mathbf{z}_{c}=\Phi_{\theta_{SE}}(\mathbf{x})$, and $\mathbf{I}+\mathbf{A}=\mathrm{d}_{\mathbf{z}_{c}}\!\Phi_{\theta_{AD}}(\mathbf{z}_{c},\mathbf{u}_{t},\mathbf{d}_{t})$ with $\|A\|\leq\rho$. Then
$\mathrm{d}\Phi_{\theta}(\mathbf{x},\mathbf{u}_{t},\mathbf{d}_{t})=(\mathbf{I}+\mathbf{A})\,\mathrm{d}\Phi_{\theta_{SE}}(\mathbf{x}).$
@@
$\mathrm{d}\Phi_{\theta}^{\top}\,\mathbf{J}\,\mathrm{d}\Phi_{\theta}=\mathrm{d}\Phi_{\theta_{SE}}^{\top}\,(\mathbf{I}+\mathbf{A})^{\top}\,\mathbf{J}\,(\mathbf{I}+\mathbf{A})\,\mathrm{d}\Phi_{\theta_{SE}}.$
Expanding $(\mathbf{I}+\mathbf{A})^{\top}\mathbf{J}(\mathbf{I}+\mathbf{A})=\mathbf{J}+\mathbf{A}^{\top}\mathbf{J}+\mathbf{J}\mathbf{A}+\mathbf{A}^{\top}\mathbf{J}\mathbf{A}=\mathbf{J}+O(\|\mathbf{A}\|)$, we substitute $\mathrm{d}\Phi_{\theta_{SE}}^{\top}\,\mathbf{J}\,\mathrm{d}\Phi_{\theta_{SE}}=\mathbf{J}$:
$\mathrm{d}\Phi_{\theta}^{\top}\,\mathbf{J}\,\mathrm{d}\Phi_{\theta}=\mathbf{J}+O(\|A\|)=\mathbf{J}+O(\rho).$
In operator norm, $\|\mathrm{d}\Phi_{\theta}^{\top}\,\mathbf{J}\,\mathrm{d}\Phi_{\theta}-\mathbf{J}\|\leq C\,\rho$. Since $\mathrm{d}\Phi_{\theta_{SE}}$ is volume-preserving, $\det(\mathrm{d}\Phi_{\theta})=\det(\mathbf{I}+\mathbf{A})\odot\mathbf{1}=\mathbf{1}+O(\rho)$. ∎
@@
$F_{\theta_{AD}}(\mathbf{z}_{c},\mathbf{u}_{t},\mathbf{d}_{t})=\;\alpha\,\mathrm{CrossAttn}(\mathbf{z}_{c},\mathbf{u}_{t})\;-\;\gamma\,(\mathbf{d}_{t}^{\top}\mathbf{z}_{c}),$
where $\mathbf{d}_{t}^{\top}\mathbf{z}_{c}$ indicates some parametric dissipator. If $\mathrm{CrossAttn}$ has partial derivative in $\mathbf{z}_{c}$ normed by $\|\mathbf{u}_{t}\|$, and $\mathbf{d}_{t}^{\top}\mathbf{z}_{c}$ is linear in $\|\mathbf{d}_{t}\|$, then
@@
We choose not to explicitly bound this perturbation during training to allow for MetaSym to model extremely dissipative systems. In Appendix D.2, we provide the perturbation bound $C\rho$ empirically for systems undergoing extreme dissipation and control. We observe that this bound is <1<1 even under these extreme settings.